what's  dz $cz2+42+13 , Where C is the contour  || 2 – 3i || = 3.
Answer #1

Here basically we use the cauchy residue formula for integrals, using the pole.

We have the contour 112-3;}|= 3 Now given integral dz Ic 2²+42 +13 ļ Thus 22+42 +13= (2-(-2+ 3i) (2 -(-2-3i)) function HenceThus the function has only z²+42 +13 one pole al z=-2+b; in the contour 112-3ill=3. As we know. For he heide the The Cauchy R톱 19. - 기 - -1+3j + +3 ul = - - +3) - (-3) -\ul = (11) -1 !G+z-42 = 2ni lim - + 러2+3 2x; lim (+) -)

Thus we are done.